OpenStudy (warriorz13):

-200 + 200 + 4(–200) + (–300)

OpenStudy (yanasidlinskiy):

Do you know what Distributive Property is??

OpenStudy (anonymous):

-200 + 200 = 0 4(-200) = -800 -800 + (-300)= -1100

OpenStudy (warriorz13):

No???? It is just telling me to find out the answer to that equation

OpenStudy (anonymous):

@chrisvillasmil i s right.